OpenStudy (anonymous):

prove geometrically, if the distances of a point from two fixed points are in constant ratio, and the ratio is not equal to 1, then the locus of the variable point will be a circle

OpenStudy (anonymous):

This is in two dimeensional space right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

I got it.

OpenStudy (anonymous):

Isn't it supposed to be an ellipse?

OpenStudy (anonymous):

i dont think that

OpenStudy (anonymous):

in ellipse, the sum of distances of any point on it from the foci are constant

OpenStudy (anonymous):

in fact, if u take 2 fixed points and write the equation for any arbitrary point satisfying the property, u will get an equation of a circle

OpenStudy (anonymous):

Suppose the fixed points are A, at (x_1,y_1), and B, at (x_2, y_2). We're given that the distance of point P, at (x,y), from points A and B is in a fixed ratio, i.e. AP/BP = k, where k is not 1. We can say then that \[AP=kBP\] AP's length is \[\sqrt{(x-x_1)^2+(y-y_1)^2}\] and BP's length is \[\sqrt{(x-x_2)^2+(y-y_2)^2}\] So we have \begin{eqnarray*} \sqrt{(x-x_1)^2+(y-y_1)^2}&=&k\sqrt{(x-x_2)^2+(y-y_2)^2} \\ (x-x_1)^2+(y-y_1)^2&=&k^2((x-x_2)^2+(y-y_2)^2) \\ x^2-2xx_1+x_1^2+y_2-2yy_1+y_1^2&=&k^2(x^2-2xx_2+x_2^2+y^2-2yy_2+y_2^2) \\ x^2-2xx_1+x_1^2+y_2-2yy_1+y_1^2&=&k^2x^2-2k^2xx_2+k^2x_2^2+k^2y^2-2k^2yy_2+k^2y_2^2 \end{eqnarray*} to be continued

OpenStudy (anonymous):

i solved by that method, but i wanna know the synthetic geometrical proof.

OpenStudy (anonymous):

so you know the answer by this method and you want a construction diagram?

OpenStudy (anonymous):

yes, exactly. can u help me, please?

OpenStudy (anonymous):

constructing this would be a bit difficult seeing as the radius of the resulting circle is quite a complicated expression, and so are the coordinates of the centre of the circle.

OpenStudy (anonymous):

ok, i will b happy if anyone answers this. thanx 4 ur co-operation

OpenStudy (anonymous):

if you want I can finish my algebraic argument. It has a subtle trick at a certain step.

OpenStudy (anonymous):

certainly, you can do. i can verify my process

OpenStudy (anonymous):

Part 2: \[(1-k^2)x^2-2x(x_1-k^2x_2)+(x_1^2-k^2x_2^2)+(1-k^2)y_2^2-2y(y_1-k^2y_2)+(y_1^2-k^2y_2^2)=0\] \[x^2-\frac{2(x_1-k^2x_2)}{1-k^2}+\frac{x_1^2-k^2x_2^2}{1-k^2}+y^2-\frac{2(y_1-k^2y_2)}{1-k^2}+\frac{y_1^2-k^2y_2^2}{1-k^2}=0\] We consider \[x^2-\frac{2(x_1-k^2x_2)}{1-k^2}+\frac{x_1^2-k^2x_2^2}{1-k^2}\] only; the other three terms would be transformed in exactly the same way. Completing the square of the term, we get: \[x^2-\frac{2(x_1-k^2x_2)}{1-k^2}+\Big(\frac{x_1-k^2x_2}{1-k^2}\Big)^2+\frac{x_1^2-k^2x_2^2}{1-k^2}-\Big(\frac{x_1-k^2x_2}{1-k^2}\Big)^2\] \[=\Big(x-\frac{x_1-k^2x_2}{1-k^2}\Big)^2+\Big(\frac{x_1^2-k^2x_2^2}{1-k^2}-\Big(\frac{x_1-k^2x_2}{1-k^2}\Big)^2\Big)\] Combining this with the other one, we get \[(x-a)^2+(y-b)^2=r^2\] where \begin{eqnarray*} a&=&\frac{x_1-k^2x_2}{1-k^2} \\ b&=&\frac{y_1-k^2y_2}{1-k^2} \\ r^2&=&\Big(\frac{x_1^2-k^2x_2^2}{k^2-1}-\Big(\frac{x_1-k^2x_2}{k^2-1}\Big)^2\Big)+\Big(\frac{y_1^2-k^2y_2^2}{k^2-1}-\Big(\frac{y_1-k^2y_2}{k^2-1}\Big)^2\Big) \end{eqnarray*} Notice also that if k=1, the denominator would vanish, which would make the whole argument above fail.